### v2

# On eigenvalues of the Landau Hamiltonian with a periodic electric potential

Abstract
We consider the Landau Hamiltonian $\widehat{H}_{B}+V$ on $L^{2}({\mathbb{R}}^{2})$ with a periodic electric potential $V$. For every $m\in{\mathbb{N}}$ we prove that there exist nonconstant periodic electric potentials $V\in C^{\infty}({\mathbb{R}}^{2};{\mathbb{R}})$ with zero mean values that analytically depend on a small parameter $\varepsilon\in{\mathbb{R}}$ such that the Landau level $(2m+1)B$ is an eigenvalue of the Hamiltonian (of infinite multiplicity) where $B>0$ is a strength of a homogeneous magnetic field.
Udmurt Federal Research Center of Ural Branch of the Russian Academy of Sciences, Izhevsk, Russia
danilov@udman.ru
Keywords: Landau Hamiltonian, periodic electric potential, spectrum.
MSC 35P05
Introduction
We consider the Landau Hamiltonian
$\widehat{H}_{B}+V=\left(-i\,\frac{\partial}{\partial x_{1}}\right)^{2}+\left(-i\,\frac{\partial}{\partial x_{2}}-Bx_{1}\right)^{2}+V$ (0.1)
acting on $L^{2}({\mathbb{R}}^{2})$ where $V$ is a periodic electric potential $V\in L^{2}_{\mathrm{loc}}({\mathbb{R}}^{2};{\mathbb{R}})$ and $B>0$ is the strength of a homogeneous magnetic field. The coordanates in ${\mathbb{R}}^{2}$ are defined with respect to some orthonomal basis $e_{1},e_{2}$. For the zero electric potential $V\equiv 0$ the spectrum of the Landau Hamiltonian
$\widehat{H}_{B}=\left(-i\,\frac{\partial}{\partial x_{1}}\right)^{2}+\left(-i\,\frac{\partial}{\partial x_{2}}-Bx_{1}\right)^{2}$
consists of the eigenvalues $\lambda=(2m+1)B$, $m\in{\mathbb{Z}}_{+}\doteq{\mathbb{N}}\cup\{0\}$, of infinite multiplicity (Landau levels). Until recently, for a rational magnetic flux, it was not known whether eigenvalues can exist in the spectrum of operator (0.1) for nonconstant periodic electric potentials $V\in C^{\infty}({\mathbb{R}}^{2};{\mathbb{R}})$ [1,2,3] (in [2] this problem was stated for nonconstant periodic potentials $V\in C({\mathbb{R}}^{2};{\mathbb{R}})$). In [4], in the case of a rational magnetic flux, it was proved that there are nonconstant periodic electric potentials $V\in C^{\infty}({\mathbb{R}}^{2};{\mathbb{R}})$ with zero mean values such that the operator (0.1) has an eigenvalue at the second Landau level $3B$. In this article for any $m\in{\mathbb{N}}$ we give examples of families of periodic electric potentials $V(x_{1},x_{2})=v(x_{2})$ such that the operator $\widehat{H}_{B}+V$ has an eigenvalue at the $(m+1)$-th Landau level $(2m+1)B$; the nonconstant $T$-periodic functions $v(\cdot)$ with a given period $T>0$ and zero mean values analytically depend on a small parameter $\varepsilon\in{\mathbb{R}}$, the strength of a magnetic field $B$ also depends on the parameter $\varepsilon$. These examples imply that for any $m\in{\mathbb{N}}$ there are nonconstant periodic electric potentials $V\in C^{\infty}({\mathbb{R}}^{2};{\mathbb{R}})$ with zero mean values such that the magnetic flux is rational and the operator $\widehat{H}_{B}+V$ has the eigenvalue $\lambda=(2m+1)B$.
0.1. Notation. Let $E^{1}$ and $E^{2}$ be basis vectors of the non-degenerate lattice $\Lambda=\{N_{1}E^{1}+N_{2}E^{2}:N_{1},N_{2}\in{\mathbb{Z}}\}$; let $E^{l}_{j}=(E^{l},e_{j})$, $l,j=1,2$ (the inner product and the length of vectors from ${\mathbb{R}}^{2}$ are denoted by $|\cdot|$ and $(\cdot,\cdot)$, respectively). By $K=\{\xi_{1}E^{1}+\xi_{2}E^{2}:0\leqslant\xi_{j}\leqslant 1,j=1,2\}$ we denote the unit cell of the lattice $\Lambda$ of area $v(K)$ corresponding to the basis vectors $E^{1},E^{2}$ where $v(\cdot)$ is the Lebesgue measure on ${\mathbb{R}}^{2}$, and let $\eta=(2\pi)^{-1}Bv(K)$ be the magnetic flux (defined for the unit cell $K$). We use the notation $\Lambda^{*}$ for the reciprocal lattice in ${\mathbb{R}}^{2}$ with basis vectors $E^{1}_{*},E^{2}_{*}\in{\mathbb{R}}^{2}$ for which $(E^{\,\mu},E^{\,\nu}_{*})=\delta_{\mu\nu}$, ${\mu},{\nu}=1,2$, where $\delta_{\mu\nu}$ is the Kronecker delta.
For $\mathcal{B}$$\mathbb{R}$} or $\mathcal{B}$$\mathbb{C}$}, let $L^{p}_{\Lambda}({\mathbb{R}}^{2};{\mathcal{B}})$, $p\in[1,+\infty]$, $C^{n}_{\Lambda}({\mathbb{R}}^{2};{\mathcal{B}})$, $n\in{\mathbb{Z}}_{+}$, $C^{\infty}_{\Lambda}({\mathbb{R}}^{2};{\mathcal{B}})$ and $H^{s}_{\Lambda}({\mathbb{R}}^{2};{\mathcal{B}})$, $s\geqslant 0$, be the spaces of functions $W:{\mathbb{R}}^{2}\to{\mathcal{B}}$ from $L^{p}_{\mathrm{loc}}({\mathbb{R}}^{2};{\mathcal{B}})$, $C^{n}({\mathbb{R}}^{2};{\mathcal{B}})$, $C^{\infty}({\mathbb{R}}^{2};{\mathcal{B}})$ and from Sobolev spaces $H^{s}_{\mathrm{loc}}({\mathbb{R}}^{2};{\mathcal{B}})$, respectively, which are periodic with a period lattice $\Lambda$ ($C^{0}_{\Lambda}({\mathbb{R}}^{2};{\mathcal{B}})=C_{\Lambda}({\mathbb{R}}^{2};{\mathbb{C}})$ and $H^{0}_{\Lambda}({\mathbb{R}}^{2};{\mathcal{B}})=L^{2}_{\Lambda}({\mathbb{R}}^{2};{\mathcal{B}})$). For functions $W\in L^{p}_{\Lambda}({\mathbb{R}}^{2};{\mathbb{C}})$ and $W\in C_{\Lambda}({\mathbb{R}}^{2};{\mathbb{C}})$ we define the norms $\|W\|_{L^{p}_{\Lambda}}\doteq\|W(\cdot\big|_{K})\|_{L^{p}(K)}$ and $\|W\|_{C_{\Lambda}}\doteq\|W(\cdot\big|_{K})\|_{C(K)}$.
Let $W_{Y}$ denote the Fourier coefficients of functions $W\in L^{1}_{\Lambda}({\mathbb{R}}^{2};{\mathbb{C}})$:
$W_{Y}=(v(K))^{-1}\int_{K}W(x)\,e^{-i(Y,x)}\,dx,\quad Y\in 2\pi\Lambda^{*};$
$W_{0}$ is the mean value of the function $W$ (where $0\in 2\pi\Lambda^{*}$ is the zero vector).
By $L^{2}_{\Lambda(T)}({\mathbb{R}};{\mathcal{B}})$, $C_{\Lambda(T)}({\mathbb{R}};{\mathcal{B}})$, and $C^{\infty}_{\Lambda(T)}({\mathbb{R}};{\mathcal{B}})$, where $\Lambda(T)=T\mathbb{Z}$, $T>0$, and $\mathcal{B}$$\mathbb{R}$} or $\mathcal{B}$$\mathbb{C}$}, we denote the spaces of $T$-periodic functions $w:{\mathbb{R}}\to{\mathcal{B}}$ from the spaces $L^{2}_{\mathrm{loc}}({\mathbb{R}};{\mathcal{B}})$, $C({\mathbb{R}};{\mathcal{B}})$, and $C^{\infty}({\mathbb{R}};{\mathcal{B}})$, respectively. For functions $w\in L^{2}_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})$ and $w\in C_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})$ we write $\|w\|_{L^{2}_{\Lambda(T)}}\doteq\|w(\cdot\big|_{[0,T]})\|_{L^{2}[0,T]}$ and $\|w\|_{C_{\Lambda(T)}}\doteq\|w(\cdot|_{[0,T]})\|_{C[0,T]}$. If $W(x_{1},x_{2})=w(x_{2})$ where $w\in L^{2}_{\Lambda(T)}({\mathbb{R}};{\mathbb{R}})$, then for the function $W$ one can choose a non-degenerate period lattice $\Lambda\subset{\mathbb{R}}^{2}$ with basis vectors $E^{1}=\alpha e_{1}$ and $E^{2}=Te_{2}$ where $\alpha$ is an arbitrary number. Then $W_{Y}=0$ if $Y_{1}\neq 0$, and
$W_{Y}=w_{Y_{2}}\doteq T^{-1}\int_{\,0}^{\,T}w(t)\,e^{-iY_{2}t}\,dt$
if $Y_{1}=0$, where $Y_{j}=(Y,e_{j})$, $j=1,2$, and $Y_{2}=2\pi T^{-1}n$, $n\in{\mathbb{Z}}$.
0.2. Some known results about the spectrum of operator (0.1). In [5], the Landau Hamiltonian $\widehat{H}_{B}+V$ was considered for perodic point potentials $V$. It was proved that there are eigenvalues (of infinite multiplicity) at the Landau levels $\lambda=(2m+1)B$, $m\in{\mathbb{Z}}_{+}$, if the period lattice $\Lambda$ is monatomic and $1<\eta\in{\mathbb{Q}}$ (and the spectrum also has an absolutely continuous component).
[…]
It was shown in [2] that, for any period lattice $\Lambda\subset{\mathbb{R}}^{2}$, in the Banach space $(C_{\Lambda}({\mathbb{R}}^{2};{\mathbb{R}}),\|\cdot\|_{C_{\Lambda}})$, there exits a dense $G_{\delta}$-set [^2] The complement of the set $\mathcal{O}$} is a set of first Baire category. $\mathcal{O}$} such that, for any potential $V\in{\mathcal{O}}$ and for any homogeneous magnetic field with the flux $0<\eta\in{\mathbb{Q}}$, the spectrum of operator (0.1) is absolutely continuous. A similar result for potentials $V\in L^{p}_{\Lambda}({\mathbb{R}}^{2};{\mathbb{R}})$, $p>1$, was obtained in [9], and for potentials from the spaces $L^{\infty}_{\Lambda}({\mathbb{R}}^{2};{\mathbb{R}})$, $C^{n}_{\Lambda}({\mathbb{R}}^{2};{\mathbb{R}})$, $n\in{\mathbb{N}}$, and $H^{s}_{\Lambda}({\mathbb{R}}^{2};{\mathbb{R}})$, $s>0$, similar results were obtained in [10]. In [10], the absolute continuity of the spectrum of operator (0.1) was also proved for potentials $C^{n}_{\Lambda}({\mathbb{R}}^{2};{\mathbb{R}})$, $n\in{\mathbb{N}}$, if $0<\eta\in{\mathbb{Q}}$ and
[…]
(for $n=0$ also see [11]).
[…]
If $V(x_{1},x_{2})=v(x_{2})$ where $v\in L^{2}_{\Lambda(T)}({\mathbb{R}};{\mathbb{R}})$, $T>0$, then one can choose the period lattice $\Lambda$ for the potential $V$ with basis vectors $E^{1}=2\pi T^{-1}B^{-1}e_{1}$ and $E^{2}=Te_{2}$ for which $\eta=1$. Therefore, the following theorem is a direct consequence of Theorem 1 and Theorem 2.
[…]
0.3. Definition of functions ${\mathcal{W}}^{(j)}$ and the main theorem. Let $C_{b}({\mathbb{R}}^{2};{\mathbb{R}})$ be the space of continuous bounded functions $W:{\mathbb{R}}^{2}\to{\mathbb{R}}$, and let $C_{b}^{\infty}({\mathbb{R}}^{2};{\mathbb{R}})$ be the space of functions $W\in C^{\infty}({\mathbb{R}}^{2};{\mathbb{R}})\cap C_{b}({\mathbb{R}}^{2};{\mathbb{R}})$ such that all partial derivatives ${\partial}^{n_{1}+n_{2}}/{\partial}x_{1}^{n_{1}}{\partial}x_{2}^{n_{2}}$, $n_{1},n_{2}\in{\mathbb{Z}}_{+}$, belong to the space $C_{b}({\mathbb{R}}^{2};{\mathbb{R}})$. By ${\mathfrak{L}}_{m}(B)$, $m\in{\mathbb{N}}$, we denote the set of ordered sets $\{{\mathcal{W}}^{(j)}\}$ of functions ${\mathcal{W}}^{(j)}\in C_{b}^{\infty}({\mathbb{R}}^{2};{\mathbb{R}})$, $j=0,\dots,m-1$, provided that the following conditions are fulfilled:
(a) ${\mathcal{W}}^{(j)}(x)<0$ for all $x\in{\mathbb{R}}^{2}$,
(a1) $1/{\mathcal{W}}^{(j)}(\cdot)\in C_{b}({\mathbb{R}}^{2};{\mathbb{R}})$ (then $1/{\mathcal{W}}^{(j)}(\cdot)\in C_{b}^{\infty}({\mathbb{R}}^{2};{\mathbb{R}})$),
(b) for all $j=1,\dots,m-1$
${\mathcal{W}}^{(j)}={\mathcal{W}}^{(0)}+2Bj-\Delta\ln\,(|{\mathcal{W}}^{(0)}|^{j}\,|{\mathcal{W}}^{(1)}|^{j-1}\dots|{\mathcal{W}}^{(j-1)}|),$
(c) ${\mathcal{W}}^{(m)}\doteq{\mathcal{W}}^{(0)}+2Bm-\Delta\ln\,(|{\mathcal{W}}^{(0)}|^{m}\,|{\mathcal{W}}^{(1)}|^{m-1}\dots|{\mathcal{W}}^{(m-1)}|)\equiv 0$.
the following theorem is the mail result of this article.
[…]
For non-degenerate lattice $\Lambda\subset{\mathbb{R}}^{2}$ let ${\mathfrak{L}}_{m}^{\Lambda}(B)$, $m\in{\mathbb{N}}$, be the set of ordered sets $\{{\mathcal{W}}^{(j)}\}\in{\mathfrak{L}}_{m}(B)$ such that $\{{\mathcal{W}}^{(j)}\}\in C_{\Lambda}^{\infty}({\mathbb{R}}^{2};{\mathbb{R}})$ for all $j=0,\dots,m-1$.
If $\{{\mathcal{W}}^{(j)}\}\in{\mathfrak{L}}_{m}^{\Lambda}(B)$, $m\in{\mathbb{N}}$, it follows from the condition (c) that ${\mathcal{W}}^{(0)}_{0}=-2Bm$. Then the condition (b) implies that the functions ${\mathcal{W}}^{(j)}$, $j=0,\dots,m-1$, have mean values ${\mathcal{W}}^{(j)}_{0}=-2B(m-j)$.
[…]
Theorem 5 is a direct consequence of Theorem 4. A particular case of Theorem 5 was proved in [4] for $\{{\mathcal{W}}^{(j)}\}\in{\mathfrak{L}}_{1}^{\Lambda}(B)$ and $\eta\in{\mathbb{Q}}$.
0.4. Existence of functions ${\mathcal{W}}^{(j)}$. For functions ${\mathcal{W}}^{(j)}\in C_{b}^{\infty}({\mathbb{R}}^{2};{\mathbb{R}})$, $j=0,\dots,m-1$, conditions (a) and (a1) hold iff
${\mathcal{W}}^{(j-1)}=-2B(m-j+1)e^{u_{j}}$ (0.2)
where $u_{j}\in C_{b}^{\infty}({\mathbb{R}}^{2};{\mathbb{R}})$, $j=1,\dots,m$. Conditions (b) and (c) are fulfilled iff
$\Delta u_{1}=2B(e^{u_{1}}-1)$ (0.3)
for $m=1$, and
$\Delta u_{1}=2Bm(e^{u_{1}}-1)-2B(m-1)(e^{u_{2}}-1),$
$j\Delta u_{1}-(j-1)\Delta u_{2}-\dots-\Delta u_{j}=2Bm(e^{u_{1}}-1)-2B(m-j)(e^{u_{j+1}}-1),\quad j=2,\dots,m-1,$
$m\Delta u_{1}-(m-1)\Delta u_{2}-\dots-\Delta u_{m}=2Bm(e^{u_{1}}-1)$ (0.4)
for $m\geqslant 2$. Furthermore, the inclusions $\{{\mathcal{W}}^{(j)}\}\in{\mathfrak{L}}_{m}^{\Lambda}(B)$, $m\in{\mathbb{N}}$, hold iff equality (0.2) and equations (0.3), (0.4) hold for functions $u_{j}\in C_{\Lambda}^{\infty}({\mathbb{R}}^{2};{\mathbb{R}})$, $j=1,\dots,m$.
If $m=1$ and $u_{1}(x_{1},x_{2})=u(x_{2})$, then equation (0.3) becomes an ordinary differential equation
$\frac{d^{2}u}{dt^{2}}=2B(e^{u}-1),\quad t\in{\mathbb{R}}.$ (0.5)
All solutions of equation (0.5) are analytic and periodic.
Lemma 1. Let $u(\alpha;\cdot)$, $\alpha>0$, be a solution of equation $\mathrm{(}0.5)$} such that $u(\alpha;0)=0$, $\frac{du(\alpha;t)}{dt}\big|_{t=0}=\alpha$, and let $T_{\alpha}>0$ be the minimal period of the function $u(\alpha;\cdot)$. Then the function $(0,+\infty)\ni\alpha\mapsto T_{\alpha}$ is continuous, $T_{\alpha}\to 2\pi(2B)^{-1/2}$ as $\alpha\to+0$ and $T_{\alpha}\to+\infty$ as $\alpha\to+\infty$.
The following theorem is a consequence of Theorem 5.
[…]
If we subtract the (j1j-1)-th equation of the system (4) from the $j$-th equation, $j=2,\dots,m$, we get the system of equations
$\Delta u_{1}=2Bm(e^{u_{1}}-1)-2B(m-1)(e^{u_{2}}-1),$
$\Delta u_{1}-\Delta u_{2}-\dots-\Delta u_{j}=2B(m-j+1)(e^{u_{j}}-1)-2B(m-j)(e^{u_{j+1}}-1),\quad j=2,\dots,m-1,$
$\Delta u_{1}-\Delta u_{2}-\dots-\Delta u_{m}=2B(e^{u_{m}}-1).$ (0.6)
If we do the same operations for the system (0.6), we get the system of equations
$\Delta u_{j}=B\sum\limits_{\mu\,=\,1}^{m}C^{(m)}_{j\mu}(e^{u_{\mu}}-1),\quad j=1,\dots,m,$ (0.7)
which is equivalent to the equation (0.3) for $m=1$ and is equivalent to the system of equations (0.4) for $m\geqslant 2$, where $C^{(m)}_{j\mu}$ are elements of a ($m\times m$)-matrix $\widehat{C}^{(m)}$, $m\in{\mathbb{N}}$; $C^{(1)}_{11}=2$ if $m=1$, and for $m\geqslant 2$ the following elements of the matrix $\widehat{C}^{(m)}$ are chosen: $C^{(m)}_{11}=2m$, $C^{(m)}_{jj}=4(m-j+1)$ if $j=2,\dots,m$, $C^{(m)}_{j,j-1}=-2(m-j+2)$ if $j=2,\dots,m$, $C^{(m)}_{j,j+1}=-2(m-j)$ if $j=1,\dots,m-1$, and $C^{(m)}_{j\mu}=0$ otherwise.
Lemma 2. All eigenvalues ${\lambda}^{(m)}_{j}$, $j=1,\dots,m$, of the matrix $\widehat{C}^{(m)}$ are real and $0<{\lambda}^{(m)}_{1}<\dots<{\lambda}^{(m)}_{m}$. Moreover, for any eigenvector [^3] The sign “T” represents the transposition of a row vector. $a=(a_{1},\dots,a_{m})^{T}$, the inequality $a_{1}\neq 0$ holds.
Let $B_{0}>0$. For a given $m\in{\mathbb{N}}$, let denote by ${\lambda}^{(m)}$ any eigenvalue of the matrix $\widehat{C}^{(m)}$ such that the following condition is fulfilled:
$n^{2}{\lambda}^{(m)}\neq{\lambda}^{(m)}_{j}\ \ for\ all\ \ n\in{\mathbb{N}}\backslash\{1\}\ \ and\ all\ \ {\lambda}^{(m)}_{j}>{\lambda}^{(m)}.$ (0.8)
Let $\omega\doteq\sqrt{B_{0}{\lambda}^{(m)}}$. By $a=(a_{1},\dots,a_{m})^{T}\in{\mathbb{R}}^{m}$ we denote the eigenvector of the matrix $\widehat{C}^{(m)}$ with the eigenvalue ${\lambda}^{(m)}$, $\|a\|_{{\mathbb{R}}^{m}}=\bigl(\,\sum a_{j}^{2}\,\bigr)^{1/2}=1$ and $a_{1}>0$. Let $T=2\pi{\omega}^{-1}$, $\Lambda(T)=T{\mathbb{Z}}$, and ${\Lambda}^{*}=T^{-1}{\mathbb{Z}}=(2\pi)^{-1}{\omega}{\mathbb{Z}}$.
Let $L^{2,\,even}_{\Lambda(T)}({\mathbb{R}};{\mathcal{B}})$, $C^{\,even}_{\Lambda(T)}({\mathbb{R}};{\mathcal{B}})$, and $C^{\infty,\,even}_{\Lambda(T)}({\mathbb{R}};{\mathcal{B}})$, where $\mathcal{B}$$\mathbb{R}$} or $\mathcal{B}$$\mathbb{C}$}, be the subspaces of even T-periodic functions from the spaces $L^{2}_{\Lambda(T)}({\mathbb{R}};{\mathcal{B}})$, $C_{\Lambda(T)}({\mathbb{R}};{\mathcal{B}})$, and $C^{\infty}_{\Lambda(T)}({\mathbb{R}};{\mathcal{B}})$, respectively.
If $u_{j}(x_{1},x_{2})=v_{j}(x_{2})$ where $v_{j}(\cdot)\in C^{\infty}_{\Lambda(T)}({\mathbb{R}};{\mathbb{R}})$, then the functions $u_{j}$ satisfy system (0.7) iff the functions $v_{j}$ satisfy the system of ordinary differential equations
[…]
For $B=B_{0}$, the T-periodic functions $\widetilde{v}_{j}=2\varepsilon a_{j}\cos\,{\omega}t$, $j=1,\dots,m$, $\varepsilon\in{\mathbb{R}}$, are solutions of the linearized system (for system (0.9))
[…]
In the following theorem the functions $\widetilde{v}_{j}$ are the leader terms in expansion in degrees of a small parameter $\varepsilon\in{\mathbb{R}}$ of solutions $v_{j}$ of nonlinear system (0.9). The strength of a homogeneous magnetic field $B>0$ also depends on the parameter $\varepsilon$.
Theorem 7. Let $B_{0}>0$ and let $m\in{\mathbb{N}}$. Suppose that the condition $\mathrm{(}0.8)$ for the eigenvalue ${\lambda}^{(m)}$ of the matrix $\widehat{C}^{(m)}$ is fulfilled. Then there exist a number $r>0$ and analytic functions444the number $r>0$ and the function ${\tau}(\cdot)$ do not depend on $B_{0}$. $(-r,r)\ni\varepsilon\mapsto{\tau}(\varepsilon)\in{\mathbb{R}}$ and $(-r,r)\ni\varepsilon\mapsto{\mathfrak{w}}_{j}(\varepsilon;\cdot)\in(C_{\Lambda(T)}({\mathbb{R}};{\mathbb{R}}),\|\cdot\|_{C_{\Lambda(T)}})$, $j=1,\dots,m$, such that, for all $\varepsilon\in(-r,r)$ the functions ${\mathfrak{w}}_{j}(\varepsilon;\cdot)$ belong to the subspace $C^{\infty,\,even}_{\Lambda(T)}({\mathbb{R}};{\mathbb{R}})$, $1+{\varepsilon}^{2}{\tau}(\varepsilon)>0$, and the functions
${\mathbb{R}}\ni t\mapsto v_{j}(t)=v_{j}(\varepsilon;t)\doteq 2\varepsilon a_{j}\cos\,{\omega}t+{\varepsilon}^{2}{\mathfrak{w}}_{j}(\varepsilon;t),\quad j=1,\dots,m,$ (0.10)
are solutions of the system
[…]
The following theorem is a consequence of Theorem 5, Theorem 7, and equality (0.2).
[…]
In Section 1 we prove Theorem 4. Some properties of solutions of equation (5) are considered in Section 2 where Lemma 1 is proved. Statements about eigenvalues and eigenvectors of the matrices $\widehat{C}^{(m)}$, $m\in{\mathbb{N}}$, are collected in Section 3. In Section 4 we prove Theorem 7.
1 Some auxiliary results and the proof of Theorem 4
Let $\|\cdot\|_{L^{2}}$ and $(\cdot,\cdot)_{L^{2}}$ be the norm and the inner product in the Hilbert space $L^{2}({\mathbb{R}}^{2};{\mathbb{C}})$. The inner product $(\cdot,\cdot)_{L^{2}}$ is supposed to be linear in the second argument. Define the operators
$\widehat{Z}_{\mp}=-i\,\frac{\partial}{\partial x_{1}}\pm\biggl(-i\,\frac{\partial}{\partial x_{2}}-Bx_{1}\biggr)$
acting on $L^{2}({\mathbb{R}}^{2};{\mathbb{C}})$ with domains $D(\widehat{Z}_{\mp})=\bigl\{\Phi\in L^{2}({\mathbb{R}}^{2};{\mathbb{C}})\cap H^{1}_{\mathrm{loc}}({\mathbb{R}}^{2};{\mathbb{C}}):\widehat{Z}_{\mp}\Phi\in L^{2}({\mathbb{R}}^{2};{\mathbb{C}})\bigr\}$. We have
$\widehat{H}_{B}=\widehat{Z}_{+}\widehat{Z}_{-}+B=\widehat{Z}_{-}\widehat{Z}_{+}-B.$
Let ${\mathcal{H}}^{(m)}_{B}$, $m\in{\mathbb{Z}}_{+}$, be subspaces of eigenfunctions of the operator $\widehat{H}_{B}$ with eigenvalues $\lambda=(2m+1)B$. The linear subspaces ${\mathcal{H}}^{(m)}_{B}\cap{\mathcal{S}}({\mathbb{R}}^{2};{\mathbb{C}})$ are dense in ${\mathcal{H}}^{(m)}_{B}$ (where ${\mathcal{S}}({\mathbb{R}}^{2};{\mathbb{C}})$ is the Schwartz space) and ${\mathcal{H}}^{(m)}_{B}\subset C^{\infty}({\mathbb{R}}^{2};{\mathbb{C}})$, $m\in{\mathbb{Z}}_{+}$. By $\widehat{P}^{(m)}$ we denote the orthogonal projection in $L^{2}({\mathbb{R}}^{2};{\mathbb{C}})$ onto the subspace ${\mathcal{H}}^{(m)}_{B}$.
If $\Phi\in{\mathcal{H}}^{(m)}_{B}$, $m\in{\mathbb{Z}}_{+}$, then $\widehat{Z}_{+}\Phi\in{\mathcal{H}}^{(m+1)}_{B}$ and $\|\widehat{Z}_{+}\Phi\|_{L^{2}}=\sqrt{2B(m+1)}\,\|\Phi\|_{L^{2}}$. If $\Phi\in{\mathcal{H}}^{(m)}_{B}$, $m\in{\mathbb{N}}$, then $\widehat{Z}_{-}\Phi\in{\mathcal{H}}^{(m-1)}_{B}$ and $\|\widehat{Z}_{-}\Phi\|_{L^{2}}=\sqrt{2Bm}\,\|\Phi\|_{L^{2}}$. For all $\Phi\in{\mathcal{H}}^{(0)}_{B}$, the equality $\widehat{Z}_{-}\Phi=0$ holds.
Let denote
${\mathcal{H}}^{\infty}_{B}\doteq\biggl\{\Phi\in L^{2}({\mathbb{R}}^{2};{\mathbb{C}}):\sum\limits_{m\,\in\,{\mathbb{Z}}_{+}}m^{p}\,\|\widehat{P}^{(m)}\Phi\|^{2}_{L^{2}}<+\infty\ \ for\ all\ \ p>0\biggr\};$
${\mathcal{S}}({\mathbb{R}}^{2};{\mathbb{C}})\subset{\mathcal{H}}^{\infty}_{B}\subset C^{\infty}({\mathbb{R}}^{2};{\mathbb{C}})$. For any function $\Phi\in{\mathcal{H}}^{\infty}_{B}$ there exists a sequence of functions ${\Phi}_{j}\in{\mathcal{S}}({\mathbb{R}}^{2};{\mathbb{C}})$, $j\in{\mathbb{N}}$, such that $\|\widehat{Z}_{+}^{n}(\Phi-\Phi_{j}\|_{L^{2}}\to 0$ as $j\to+\infty$ for all $n\in{\mathbb{Z}}_{+}$. The last assertion and the closedness of the operators $\widehat{Z}_{\mp}$ imply that the equalities
$\widehat{Z}_{\mp}{\mathcal{W}}\Phi=-i\,\biggl(\frac{\partial{\mathcal{W}}}{\partial x_{1}}\pm i\,\frac{\partial{\mathcal{W}}}{\partial x_{2}}\biggr)+{\mathcal{W}}\widehat{Z}_{\mp}\Phi$ (1.1)
hold for all functions $\Phi\in{\mathcal{H}}^{\infty}_{B}$ and ${\mathcal{W}}\in C^{\infty}_{b}({\mathbb{R}}^{2};{\mathbb{C}})$.
Right inverses of the operator $\widehat{Z}_{-}$ are bounded linear operators but are not uniquely defined. Let $\widehat{Z}_{-}^{-1}$ be the operator such that
$\widehat{Z}_{-}^{-1}\Phi=\sum\limits_{m\,\in\,{\mathbb{Z}}_{+}}\bigl(2B(m+1)\bigr)^{-1}\widehat{Z}_{+}\widehat{P}^{(m)}\Phi,\quad\Phi\in L^{2}({\mathbb{R}}^{2};{\mathbb{C}}).$
[…]
$\widehat{Z}_{-}^{-1}\widehat{Z}_{-}\Phi=\Phi-\widehat{P}^{(0)}\Phi.$ (1.2)
For any function $\Phi\in{\mathcal{H}}^{\infty}_{B}$, we have $\widehat{Z}_{\mp}\Phi\in{\mathcal{H}}^{\infty}_{B}$ and $\widehat{Z}_{-}^{-1}\Phi\in{\mathcal{H}}^{\infty}_{B}$.
Lemma 3. If $\Phi\in{\mathcal{H}}^{\infty}_{B}$ and ${\mathcal{W}}\in C^{\infty}_{b}({\mathbb{R}}^{2};{\mathbb{C}})$, then ${\mathcal{W}}\Phi\in{\mathcal{H}}^{\infty}_{B}$.
Proof. Let $\Phi^{\prime}\in{\mathcal{H}}^{(m)}_{B}$, $m\in{\mathbb{Z}}_{+}$, and $q\in{\mathbb{N}}$. Because
$(\Phi^{\prime},{\mathcal{W}}\Phi)_{L^{2}}=\bigl((2B)^{q}(m+1)\dots(m+q)\bigr)^{-1}(\Phi^{\prime},\widehat{Z}_{-}^{q}\widehat{Z}_{+}^{q}{\mathcal{W}}\Phi)_{L^{2}},$
it follows from (1.1) that
$\|\widehat{P}^{(m)}{\mathcal{W}}\Phi\|_{L^{2}}\leqslant C\bigl((2B)^{q}(m+1)\dots(m+q)\bigr)^{-1}$
where $C=C(B,q;{\mathcal{W}},\Phi)>0$. Hence, ${\mathcal{W}}\Phi\in{\mathcal{H}}^{\infty}_{B}$. $\square$
Let $\{{\mathcal{W}}^{(j)}\}\in{\mathfrak{L}}_{m}(B)$, $m\in{\mathbb{N}}$, and let
$\widehat{\mathcal{B}}^{(0)}=\widehat{Z}_{-}^{-1}(\widehat{Z}_{+}\widehat{Z}_{-}+{\mathcal{W}}^{(0)}).$ (1.3)
Sequentially for $j=1,\dots,m$ we define the operators
$\widehat{\mathcal{B}}^{(j)}\doteq({\mathcal{W}}^{(j-1)})^{-1}\widehat{Z}_{-}\widehat{\mathcal{B}}^{(j-1)}\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(j-1)}.$ (1.4)
All these operators can be considered as operators acting on ${\mathcal{H}}^{\infty}_{B}$.
Lemma 4. For all $j=1,\dots,m$,
$\widehat{\mathcal{B}}^{(j)}=\widehat{Z}_{+}+\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(j-1)}-i\,\biggl(\biggl(\frac{\partial}{\partial x_{1}}-i\,\frac{\partial}{\partial x_{2}}\biggr)\ln\,|{\mathcal{W}}^{(0)}\dots\mathcal{W}^{(j-1)}|\biggr),$ (1.5)
and for all $j=0,1,\dots,m$ and $\Psi\in{\mathcal{H}}_{B}^{(0)}$,
$\widehat{Z}_{-}\widehat{\mathcal{B}}^{(j)}\Psi={\mathcal{W}}^{(j)}\Psi.$ (1.6)
Proof. We have
$\widehat{\mathcal{B}}^{(1)}\doteq({\mathcal{W}}^{(0)})^{-1}\widehat{Z}_{-}\widehat{\mathcal{B}}^{(0)}\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(0)}=({\mathcal{W}}^{(0)})^{-1}(\widehat{Z}_{+}+{\mathcal{W}}^{(0)}\widehat{Z}_{-}^{-1}){\mathcal{W}}^{(0)}$
$\widehat{Z}_{+}+\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(0)}-i\,\biggl(\biggl(\frac{\partial}{\partial x_{1}}-i\,\frac{\partial}{\partial x_{2}}\biggr)\ln\,|{\mathcal{W}}^{(0)}|\biggr).$
Now, assume that (1.5) holds for some $j\in{\mathbb{N}}$, $j<m$. Then
$\widehat{\mathcal{B}}^{(j+1)}=({\mathcal{W}}^{(j)})^{-1}\widehat{Z}_{-}\widehat{\mathcal{B}}^{(j)}\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(j)}$
$({\mathcal{W}}^{(j)})^{-1}(\widehat{Z}_{+}+(2B+{\mathcal{W}}^{(j-1)}-\Delta\ln\,|{\mathcal{W}}^{(0)}\dots\mathcal{W}^{(j-1)}|)\widehat{Z}_{-}^{-1}){\mathcal{W}}^{(j)}$
$i\,\bigg(\biggl(\frac{\partial}{\partial x_{1}}-i\,\frac{\partial}{\partial x_{2}}\biggr)\ln\,|{\mathcal{W}}^{(0)}\dots\mathcal{W}^{(j-1)}|\biggr)$
$\widehat{Z}_{+}-i\,\bigg(\biggl(\frac{\partial}{\partial x_{1}}-i\,\frac{\partial}{\partial x_{2}}\biggr)\ln\,|{\mathcal{W}}^{(j)}|\biggr)+\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(j)}$
$i\,\bigg(\biggl(\frac{\partial}{\partial x_{1}}-i\,\frac{\partial}{\partial x_{2}}\biggr)\ln\,|{\mathcal{W}}^{(0)}\dots\mathcal{W}^{(j-1)}|\biggr)$
$\widehat{Z}_{+}+\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(j)}-i\,\biggl(\biggl(\frac{\partial}{\partial x_{1}}-i\,\frac{\partial}{\partial x_{2}}\biggr)\ln\,|{\mathcal{W}}^{(0)}\dots\mathcal{W}^{(j-1)}{\mathcal{W}}^{(j)}|\biggr).$
We have shown that (1.5) is fulfilled for $j+1$, therefore (1.5) follows by induction for all $j=1,\dots,m$. By (1.3), the equality
$\widehat{Z}_{-}\widehat{\mathcal{B}}^{(0)}\Psi={\mathcal{W}}^{(0)}\Psi$
holds for all $\Psi\in{\mathcal{H}}^{(0)}_{B}$. Hence, using (1.5), for all $j=1,\dots,m$ and $\Psi\in{\mathcal{H}}^{(0)}_{B}$, we obtain
$\widehat{Z}_{-}\widehat{\mathcal{B}}^{(j)}\Psi=(2B+{\mathcal{W}}^{(j-1)}-\Delta\ln\,|{\mathcal{W}}^{(0)}\dots\mathcal{W}^{(j-1)}|)\Psi={\mathcal{W}}^{(j)}\Psi.$
Lemma 4 is proved.
THE PROOF OF THEOREM 4. It follows from (1.6) that, for all $j=1,\dots,m$ and for any function $\Psi\in{\mathcal{H}}^{(0)}_{B}$, there is a function $\Psi_{j,1}\in{\mathcal{H}}^{(0)}_{B}$ such that
$\widehat{\mathcal{B}}^{(j)}\Psi=\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(j)}\Psi+\Psi_{j,1}.$ (1.7)
For $j=1,\dots,m$, under sequential action of the operators $\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(j-1)},\dots,\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(0)}$ on left-hand side and right-hand side of equality (1.7), using (1.2), (1.3), and (1.4), we get the equations
$\widehat{\mathcal{B}}^{(0)}\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(0)}\dots\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(j-1)}\Psi$
$\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(0)}\dots\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(j-1)}\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(j)}\Psi+\sum\limits_{s\,=\,1}^{j}\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(0)}\dots\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(j-s)}\Psi_{j,s}$ (1.8)
where $\Psi_{j,s}\in{\mathcal{H}}^{(0)}_{B}$, $s=1,\dots,j$. Since ${\mathcal{W}}^{(m)}\equiv 0$, for all $\Psi\in{\mathcal{H}}^{(0)}_{B}$, we have
$\widehat{\mathcal{B}}^{(0)}\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(0)}\dots\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(m-1)}\Psi=\sum\limits_{s\,=\,1}^{m}\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(0)}\dots\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(m-s)}\Psi_{m,s}.$ (1.9)
From (1.8), (1.9) and from the equality $\widehat{\mathcal{B}}^{(0)}\Psi=\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(0)}\Psi$, it follows that for any function $\Psi\in{\mathcal{H}}^{(0)}_{B}$ we can sequentially choose functions $\Psi_{0},\dots,\Psi_{m-1}\in{\mathcal{H}}^{(0)}_{B}$ (which are uniquely determined) such that, for the function
$\Phi(\Psi)\doteq\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(0)}\dots\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(m-1)}\Psi+\sum\limits_{s\,=\,0}^{m-2}\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(0)}\dots\widehat{Z}_{-}^{-1}{\mathcal{W}}^{(m-2-s)}\Psi_{s}+\Psi_{m-1}\in{\mathcal{H}}^{(0)}_{B}$, (1.10)
the equality $\widehat{\mathcal{B}}^{(0)}(\Phi(\Psi))\equiv 0$ holds (in (1.10) the sum is absent if $m=1$). Using the equality $({\mathcal{W}}^{(m-1)})^{-1}\widehat{Z}_{-}\dots({\mathcal{W}}^{(0)})^{-1}\widehat{Z}_{-}\Phi(\Psi)=\Psi$, we obtain that $\Phi(\Psi)\not\equiv 0$ if $\Psi\not\equiv 0$. Moreover, for any finite set of linearly independent functions $\Psi\in{\mathcal{H}}^{(0)}_{B}$, the functions $\Phi(\Psi)$ also are linearly independent. By (1.3),
$(\widehat{Z}_{+}\widehat{Z}_{-}+{\mathcal{W}}^{(0)})\,\Phi(\Psi)\equiv 0$
for all non-zero functions $\Psi\in{\mathcal{H}}^{(0)}_{B}$. This concludes the proof of Theorem 4.
2 The proof of Lemma 1
Let $u_{+}(\alpha)$ and $u_{-}(\alpha)$ denote the maximum value and the minimum value of the solution $u(\alpha;\cdot)$, $\alpha>0$, of equation (0.5). Then
$e^{u_{+}(\alpha)}-1-u_{+}(\alpha)=e^{u_{-}(\alpha)}-1-u_{-}(\alpha)=(4B)^{-1}{\alpha}^{2}$
and
$\biggl(\frac{du(\alpha;t)}{dt}\biggr)^{2}={\alpha}^{2}-4B(e^{u(\alpha;t)}-1-u(\alpha;t)).$
Thus, for the (minimal) periods $T_{\alpha}$ of oscillations of the solutions $u(\alpha;\cdot)$, we obtain
$T_{\alpha}=2\int_{u_{-}(\alpha)}^{\,u_{+}(\alpha)}\frac{du}{\sqrt{{\alpha}^{2}-4B(e^{u}-1-u)}}=\alpha B^{-1}\int_{-1}^{\,1}\frac{\xi}{e^{u_{\xi}}-1}\,\frac{d\xi}{\sqrt{1-{\xi}^{2}}}$ (2.1)
where $1,1]\ni\xi\mapsto u_{\xi}\in{\mathbb{R}}$ is a monotonically increasing function such that $e^{u_{\xi}}-1-u_{\xi}=(4B)^{-1}{\alpha}^{2}{\xi}^{2}$; $u_{-1}=-(4B)^{-1}{\alpha}^{2}$, $u_{0}=0$, $u_{1}=(4B)^{-1}{\alpha}^{2}$. Equality (2.1) implies that the function $(0,+\infty)\ni\alpha\mapsto T_{\alpha}$ is continuous. If $\alpha\to+0$, then $\xi(e^{u_{\xi}}-1)^{-1}\to 2\sqrt{B}{\alpha}^{-1}$ uniformly in $\xi\in[-1,1]$. Therefore [^5] We have $T_{\alpha}=2\pi(2B)^{-1/2}\bigl(1+\frac{1}{48}\,{\alpha}^{2}B^{-1}+O({\alpha}^{4}B^{-2})\bigr)$ as $\alpha\to+0$., $T_{\alpha}\to 2\pi(2B)^{-1/2}$ as $\alpha\to+0$. On the other hand, $u_{\xi}>(4B)^{-1}{\alpha}^{2}{\xi}^{2}$ with $\xi\in[-1,0)$. Therefore, $\xi(e^{u_{\xi}}-1)^{-1}\to-\xi$ as $\alpha\to+\infty$ uniformly in $\xi\in[-1,-1/2]$, and hence [^6] Moreover, $T_{\alpha}=(1+o(1))\alpha B^{-1}$ as $\alpha\to+\infty$. $T_{\alpha}\to+\infty$ as $\alpha\to+\infty$.
3 Eigenvalues of matrices $\widehat{C}^{(m)}$, $m\in{\mathbb{N}}$
3.1. The proof of Lemma 2. We have $\widehat{C}^{(m)}=\widehat{\mathcal{E}}^{(m)}\widehat{\mathcal{D}}^{(m)}$ where $\widehat{\mathcal{E}}^{(m)}$ is a symmetric ($m\times m$)-matrix with non-zero elements ${\mathcal{E}}^{(m)}_{11}=1$ and (for $m\geqslant 2$) ${\mathcal{E}}^{(m)}_{jj}=2$ if $j=2,\dots,m$, and ${\mathcal{E}}^{(m)}_{j,\,j+1}={\mathcal{E}}^{(m)}_{j+1,\,j}=-1$ if $j=1,\dots,m-1$. The ($m\times m$)-matrix $\widehat{\mathcal{D}}^{(m)}$ is dioganal with non-zero elements ${\mathcal{D}}^{(m)}_{jj}=2(m-j+1)$, $j=1,\dots,m$.
Lemma 5. The matrix $\widehat{\mathcal{E}}^{(m)}$ is positive definite.
Proof. Let $D_{N}$, $N\in{\mathbb{N}}$, be the determinant of the symmetric ($N\times N$)-matrix ${\widehat{F}}^{(N)}$ with non-zero elements $F^{(N)}_{jj}=2$, $j=1,\dots,N$, and (for $N\geqslant 2$) $F^{(N)}_{j,\,j+1}=F^{(N)}_{j+1,\,j}=-1$, $j=1,\dots,N-1$. Since $D_{1}=2$, $D_{2}=3$, and $D_{N+2}=2D_{N+1}-D_{N}$ if $N\geqslant 1$, it follows that $D_{N}=N+1$ for all $N\in{\mathbb{N}}$. On the other hand, $\det\,\widehat{\mathcal{E}}^{(m)}=1$ if $m=1,2$, and $\det\,\widehat{\mathcal{E}}^{(m)}=D_{m-1}-D_{m-2}$ if $m\geqslant 3$. Hence $\det\,\widehat{\mathcal{E}}^{(m)}=1$ for all $m\in{\mathbb{N}}$. Thus, all minors $\det\,({\mathcal{E}}_{jk})_{j,\,k=n,\dots,N}$, $n\in\{1,\dots,N\}$, of the matrix $\widehat{\mathcal{E}}^{(m)}$ are positive and, therefore, the matrix $\widehat{\mathcal{E}}^{(m)}$ is positive definite. $\square$
Let $(\cdot,\cdot)_{{\mathbb{C}}^{m}}$ and $\|\cdot\|_{{\mathbb{C}}^{m}}$ be the inner product and the norm in ${\mathbb{C}}^{m}$; $(z^{(1)},z^{(2)})_{{\mathbb{C}}^{m}}=\sum\limits_{j\,=\,1}^{m}\overline{z}^{(1)}_{j}z^{(2)}_{j}$, $z^{(\mu)}=(z^{(\mu)}_{1},\dots,z^{(\mu)}_{m})^{T}$, $z^{(\mu)}_{j}\in{\mathbb{C}}$, $j=1,\dots,m$, $\mu=1,2$ (for real vectors we replace ${\mathbb{C}}^{m}$ with ${\mathbb{R}}^{m}$ in notation); $\widehat{I}_{m}$ is the identity ($m\times m$)-matrix.
Lemma 6. All eigenvalues of the matrix $\widehat{\mathcal{E}}^{(m)}$ are real and positive.
Proof. Let $a$ be an eigenvector of the matrix $\widehat{\mathcal{E}}^{(m)}$ with an eigenvalue $\lambda\in{\mathbb{C}}$. Then it follows from Lemma 5 that
$0<(\widehat{\mathcal{D}}^{(m)}a,\widehat{\mathcal{E}}^{(m)}\widehat{\mathcal{D}}^{(m)}a)_{{\mathbb{C}}^{m}}=(\widehat{\mathcal{D}}^{(m)}a,\widehat{C}^{(m)}a)_{{\mathbb{C}}^{m}}=\lambda\,(\widehat{\mathcal{D}}^{(m)}a,a)_{{\mathbb{C}}^{m}}.$
On the other hand, $(\widehat{\mathcal{D}}^{(m)}a,a)_{{\mathbb{C}}^{m}}>0$. Hence $\lambda>0$. $\square$
Since, for any $\lambda\in{\mathbb{C}}$, the first m1m-1 columns of the matrix $\widehat{C}^{(m)}-\lambda\widehat{I}_{m}$ are linearly independent, the inequality ${\mathrm{Rank}}\,(\widehat{C}^{(m)}-\lambda\widehat{I}_{m})\geqslant m-1$ holds and, therefore, the following lemma is fulfilled.
Lemma 7. Geometric multiplicity of eigenvalues of the matrix $\widehat{C}^{(m)}$ equals 1.
Lemma 8. Algebraic multiplicity of eigenvalues of the matrix $\widehat{C}^{(m)}$ equals 1.
Proof. Lemma 7 implies that it suffices to show that there are no associated vectors for eigenvectors of the matrix $\widehat{C}^{(m)}$. We argue by contradiction; suppose that a vector $a^{\prime}$ is an associated vector for an eigenvector $a$ with an eigenvalue $\lambda>0$. Then $\widehat{C}^{(m)}a^{\prime}=\lambda a^{\prime}+a$ and
$0=(\widehat{\mathcal{D}}^{(m)}a^{\prime},(\widehat{C}^{(m)}-\lambda\widehat{I}_{m})a)_{{\mathbb{C}}^{m}}=(\widehat{\mathcal{D}}^{(m)}a^{\prime},(\widehat{\mathcal{E}}^{(m)}-\lambda(\widehat{\mathcal{D}}^{(m)})^{-1})\widehat{\mathcal{D}}^{(m)}a)_{{\mathbb{C}}^{m}}$
$((\widehat{\mathcal{E}}^{(m)}-\lambda(\widehat{\mathcal{D}}^{(m)})^{-1})\widehat{\mathcal{D}}^{(m)}a^{\prime},\widehat{\mathcal{D}}^{(m)}a)_{{\mathbb{C}}^{m}}=\,((\widehat{C}^{(m)}-\lambda\widehat{I}_{m})a^{\prime},\widehat{\mathcal{D}}^{(m)}a)_{{\mathbb{C}}^{m}}=(a,\widehat{\mathcal{D}}^{(m)}a)_{{\mathbb{C}}^{m}}>0.$
We have arrived at a contradiction; this proves Lemma 8.
Lemma 9. If $a=(a_{1},\dots,a_{m})^{T}$ is an eigenvector of the matrix $\widehat{C}^{(m)}$, then $a_{1}\neq 0$.
Proof. It may be assumed that $m\geqslant 2$. Denote by $a$ an eigenvector of the matrix $\widehat{C}^{(m)}$ with an eigenvalue $\lambda$. Let $\widehat{\mathcal{C}}$ be the matrix which is obtained from the matrix $\widehat{C}^{(m)}-\lambda\widehat{I}_{m}$ by deletion of the first column and the last row. Suppose that $a_{1}=0$, then, for the vector $\widetilde{a}=(a_{2},\dots,a_{m})^{T}$, the vector $\widehat{\mathcal{C}}\,\widetilde{a}$ is zero. On the other hand, ${\mathrm{Rank}}\,\widehat{\mathcal{C}}=m-1$, hence $a_{2}=\dots=a_{m}=0$. Contradiction; Lemma 9 is proved.
Now, Lemma 2 follows from Lemma 6, Lemma 8, and Lemma 9.
3.2. Lemma 10 and its proof. In what follows, eigenvectors $a=(a_{1},\dots,a_{m})^{T}$ of the matrix $\widehat{C}^{(m)}$ will be chosen from the space ${\mathbb{R}}^{m}$. Let $B_{0}>0$ and let $\omega=\sqrt{B_{0}\lambda^{(m)}}$ where $\lambda^{(m)}\in\{\lambda^{(m)}_{1},\dots,\lambda^{(m)}_{m}\}$. Denote by $a=(a_{1},\dots,a_{m})^{T}\in{\mathbb{R}}^{m}$ the eigenvector of the matrix $\widehat{C}^{(m)}$ with the eigevalue $\lambda^{(m)}$ such that $\|a\|_{{\mathbb{R}}^{m}}=1$ and $a_{1}>0$. Lemma 8 implies that
$\bigl\{(\omega^{2}\widehat{I}_{m}-B_{0}\widehat{C}^{(m)})b:b\in{\mathbb{C}}^{m}\ \,and\ \,(b,a)_{{\mathbb{C}}^{m}}=0\bigr\}$
is an (m1m-1)-dimensional linear subspace of ${\mathbb{C}}^{m}$; this subspace coincides with a linear span of the eigenvectors of the matrix $\widehat{C}^{(m)}$ with eigenvalues $\lambda^{(m)}_{j}\neq\lambda^{(m)}$. Since the eigenvector $a$ does not belong to this subspace, the following lemma is valid.
Lemma 10. For any vector $f=(f_{1},\dots,f_{m})^{T}\in{\mathbb{C}}^{m}$, there exist a unique number $\tau\in{\mathbb{C}}$ and a unique vector $b=(b_{1},\dots,b_{m})^{T}\in{\mathbb{C}}^{m}$ such that $(b,a)_{{\mathbb{C}}^{m}}=0$ and
$\bigl(\omega^{2}\widehat{I}_{m}-B_{0}\widehat{C}^{(m)}\bigr)b-\tau{\omega}^{2}a\,=\,B_{0}f.$
[…]
$\tau=\sum\limits_{\mu}\tau_{\mu}f_{\mu},\quad b_{\nu}=\sum\limits_{\mu}R_{\nu\mu}f_{\mu},\quad\nu=1,\dots,m,$
and hence
$|\tau|+\sum\limits_{\nu}|b_{\nu}|\,\leqslant\,C\,\sum\limits_{\nu}|f_{\nu}|$
where $C=C(m)>0$.
4 The proof of Theorem 7
Let $T=2\pi{\omega}^{-1}$, let ${\mathbb{Y}}=\{\omega,-\omega\}$, and let ${\mathbb{Y}}^{\prime}\doteq\omega{\mathbb{Z}}\,\backslash{\mathbb{Y}}$. For Fourier coefficients of functions $v\in L^{2}_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})$, we use the notation $v_{Y}=T^{-1}\int_{0}^{T}v(t)\,e^{-iYt}\,dt$, $Y\in\omega{\mathbb{Z}}$. We write
${\mathfrak{L}}({\mathbb{Y}}^{\prime})=\bigl\{v\in L^{2,\,even}_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}}):\,v_{Y}=0\ \,for\ \,Y\in{\mathbb{Y}}\bigr\}.$
[…]
${\omega}_{*}(t)=2\cos\omega t,\quad t\in{\mathbb{R}}.$
Lemma 11. For any functions ${\mathcal{F}}_{j}\in{\mathfrak{L}}({\mathbb{Y}}^{\prime})$, $j=1,\dots,m$, there exist unique functions $G_{j}\in{\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap H^{2}_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})\subset{\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap C_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})$ such that
[…]
Moreover, $G_{j}=\sum\limits_{\mu}\widehat{S}_{j\mu}{\mathcal{F}}_{\mu}$ where $\widehat{S}_{j\mu}:({\mathfrak{L}}({\mathbb{Y}}^{\prime}),\|\cdot\|_{L^{2}_{\Lambda(T)}})\to({\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap C_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}}),\|\cdot\|_{C_{\Lambda(T)}})$ are linear bounded operators. For all ${\mathcal{F}}_{j}\in{\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap C_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})$,
$\sum\limits_{j}\|G_{j}\|_{C_{\Lambda(T)}}\leqslant CT^{-1/2}\sum\limits_{j}\|{\mathcal{F}}_{j}\|_{L^{2}_{\Lambda(T)}}\leqslant C\sum\limits_{j}\|{\mathcal{F}}_{j}\|_{C_{\Lambda(T)}}$
where $C=C(m)>0$. If ${\mathcal{F}}_{j}\in{\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap L^{2}_{\Lambda(T)}({\mathbb{R}};{\mathbb{R}})$, $j=1,\dots,m$, then $G_{j}\in{\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap C_{\Lambda(T)}({\mathbb{R}};{\mathbb{R}})$. If ${\mathcal{F}}_{j}\in{\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap C^{\infty}_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})$, $j=1,\dots,m$, then $G_{j}\in{\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap C^{\infty}_{\Lambda(T)}({\mathbb{R}};{\mathbb{R}})$.
[…]
$n^{2}{\omega}^{2}b_{j}^{(Y)}-B_{0}\sum\limits_{\mu}C^{(m)}_{j\mu}b_{\mu}^{(Y)}=B_{0}({\mathcal{F}}_{j})_{Y},$ (4.1)
and
$\sum\limits_{j}|b_{j}^{(Y)}|\,\leqslant\,C(Y)\sum\limits_{j}|({\mathcal{F}}_{j})_{Y}|$ (4.2)
where $C(Y)>0$. This implies that, for all $Y=n\omega\in{\mathbb{Y}}^{\prime}$,
$\sum\limits_{j}(1+n^{2})|b_{j}^{(Y)}|\,\leqslant\,C^{\prime}\sum\limits_{j}|({\mathcal{F}}_{j})_{Y}|$
where $C^{\prime}=C^{\prime}(m)>0$. Then, for functions
$G_{j}\doteq\sum\limits_{Y\,\in\,{\mathbb{Y}}^{\prime}}b_{j}^{(Y)}\,e^{\,iYt}\,\in\,{\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap C_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})$
we find
$\sum\limits_{j}\|G_{j}\|_{C_{\Lambda(T)}}\,\leqslant\,\sum\limits_{j}\sum\limits_{Y\,\in\,{\mathbb{Y}}^{\prime}}|(G_{j})_{Y}|$ (4.3)
$\leqslant\,\sum\limits_{j}\,\biggl(\,\sum\limits_{Y\,\in\,2\pi\Lambda^{*}(T)}(1+n^{2})^{-2}\biggr)^{1/2}\biggl(\,\sum\limits_{Y\,\in\,{\mathbb{Y}}^{\prime}}(1+n^{2})^{2}|(G_{j})_{Y}|^{2}\biggr)^{1/2}$
$\leqslant\,mC^{\prime}\biggl(\,\sum\limits_{Y\,\in\,2\pi\Lambda^{*}(T)}(1+n^{2})^{-2}\biggr)^{1/2}\biggl(\,\sum\limits_{Y\,\in\,{\mathbb{Y}}^{\prime}}\biggl(\,\sum\limits_{j}|({\mathcal{F}}_{j})_{Y}|^{2}\biggr)\biggr)^{1/2}$
$\leqslant\,C\sum\limits_{j}\,\biggl(\,\sum\limits_{Y\,\in\,{\mathbb{Y}}^{\prime}}|({\mathcal{F}}_{j})_{Y}|^{2}\biggr)^{1/2}=\,CT^{-1/2}\sum\limits_{j}\|{\mathcal{F}}_{j}\|_{L^{2}_{\Lambda(T)}}\leqslant\,\sum\limits_{j}\|{\mathcal{F}}_{j}\|_{C_{\Lambda(T)}}$
where $C=mC^{\prime}\biggl(\,\sum\limits_{Y\,\in\,2\pi\Lambda^{*}(T)}(1+n^{2})^{-2}\biggr)^{1/2}$. The functions $G_{j}$ satisfy all conditions of Lemma 11. $\square$
We use the notation $U_{r}=\{z\in{\mathbb{C}}:|z|<r\}$, $\overline{U}_{r}=\{z\in{\mathbb{C}}:|z|\leqslant r\}$, $r>0$.
Lemma 12. If, for some $r>0$, functions
$U_{r}\ni z\mapsto{\mathcal{F}}_{j}(z)\in({\mathfrak{L}}({\mathbb{Y}}^{\prime}),\|\cdot\|_{L^{2}_{\Lambda(T)}}),\quad j=1,\dots,m,$
are analytic, then the functions
$U_{r}\ni z\mapsto G_{j}(z)=\sum\limits_{\mu\,=\,1}^{m}\widehat{S}_{j\mu}{\mathcal{F}}_{\mu}(z)\in({\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap C_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}}),\|\cdot\|_{C_{\Lambda(T)}})$ (4.4)
also are analytic.
Proof. It follows from (4.1) and (4.2) that Fourier coefficients of the functions $G_{j}(z)$ are analytic functions. On the other hand, for all $R\geqslant 0$ and $z\in U_{r}$, analogously to inequality (4.3), we obtain
$\sum\limits_{j}\,\sum\limits_{|Y|\,>\,R}|(G_{j}(z))_{Y}|\,\leqslant\,mC^{\prime}\biggl(\,\sum\limits_{n\,\in\,{\mathbb{Z}}:\,|n|\omega\,>\,R}(1+n^{2})^{-2}\biggr)^{1/2}T^{-1/2}\sum\limits_{j}\|{\mathcal{F}}_{j}\|_{L^{2}_{\Lambda(T)}}.$
Therefore, for all $r_{1}\in(0,r)$, the partial sums
$\sum\limits_{Y\,\in\,2\pi\Lambda^{*}(T):\,|Y|\,\leqslant\,R}(G_{j}(z))_{Y}\,e^{\,iYt}$
converge in th space $({\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap C_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}}),\|\cdot\|_{C_{\Lambda(T)}})$ to the functions $G_{j}(z)$ as $R\to+\infty$ uniformly in $z\in\overline{U}_{r_{1}}$ and consequently functions (4.4) are analytic. $\square$
Let $B_{0}>0$. We choose a number $r>0$. Some bounds from above for the number $r$ will be given in what follows. For all $\varepsilon\in U_{r}$, let us represent complex solutions of system of equations (0.11) with $\tau=\tau(\varepsilon)\in{\mathbb{C}}$ in the form
$v_{j}=v_{j}(\varepsilon;\cdot)=\varepsilon a_{j}w_{*}+{\varepsilon}^{2}w_{j}+{\varepsilon}^{3}b_{j}w_{*},\quad j=1,\dots,m,$ (4.5)
where $a=(a_{1},\dots,a_{m})^{T}\in{\mathbb{R}}^{m}$ is the eigevector of the matrix $\widehat{C}^{(m)}$ with the eigenvalue ${\lambda}^{(m)}$, $\|a\|_{{\mathbb{R}}^{m}}=1$, $a_{1}>0$, $b=b(\varepsilon)=(b_{1},\dots,b_{m})^{T}\in{\mathbb{C}}^{m}$ and $(b,a)_{{\mathbb{C}}^{m}}=0$, $w_{j}=w_{j}(\varepsilon;\cdot)\in{\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap C^{\infty}_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})$. The last summand in (4.5) is absent if $m=1$. The (complex) functions ${\mathfrak{w}}_{j}(\varepsilon;\cdot)$ in (0.10) are chosen as ${\mathfrak{w}}_{j}(\varepsilon;\cdot)=w_{j}(\varepsilon;\cdot)+\varepsilon b_{j}w_{*}(\cdot)$, $j=1,\dots,m$.
Since $a$ is the eigenvector of the matrix $B_{0}\widehat{C}^{(m)}$ with the eigenvalue $\omega^{2}$ and $\widehat{P}_{\,{\mathbb{Y}}}w^{2}_{*}$ is the zero function, system of equations (0.11) is equivalent to systems
$\biggl(\omega^{2}b_{j}-B_{0}\sum\limits_{\mu}C^{(m)}_{j\mu}b_{\mu}-\tau\omega^{2}a_{j}\biggr)w_{*}$ (4.6)
$B_{0}\widehat{P}_{\,{\mathbb{Y}}}\biggl(\,\sum\limits_{\mu}C^{(m)}_{j\mu}a_{\mu}w_{*}w_{\mu}\biggr)+\frac{B_{0}}{6}\,\widehat{P}_{\,{\mathbb{Y}}}\biggl(\,\sum\limits_{\mu}C^{(m)}_{j\mu}a^{3}_{\mu}\biggr)w^{3}_{*}+\varepsilon B_{0}\widehat{P}_{\,{\mathbb{Y}}}{\mathcal{F}}_{j}^{(0)},\quad j=1,\dots,m,$
$\frac{d^{2}w_{j}}{dt^{2}}\,-\,B_{0}\sum\limits_{\mu}C^{(m)}_{j\mu}w_{\mu}=\frac{B_{0}}{2}\,\biggl(\,\sum\limits_{\mu}C^{(m)}_{j\mu}a^{2}_{\mu}\biggr)w^{2}_{*}+\varepsilon B_{0}\widehat{P}_{\,{\mathbb{Y}}^{\prime}}{\mathcal{F}}_{j}^{(1)},\quad j=1,\dots,m.$ (4.7)
which are derived from (0.11) under the orthogonal projections $\widehat{P}_{\,{\mathbb{Y}}}$ and $\widehat{P}_{\,{\mathbb{Y}}^{\prime}}$. The functions ${\mathcal{F}}_{j}^{(\sigma)}$, $\sigma=0,1$, can be explicitly found from (0.11) by expansion of the function ${\mathbb{C}}\ni\zeta\mapsto e^{\,\zeta}$ in power series. This functions satisfy following properties (1), (2), and (3).
(1) For all $\varepsilon,\tau,b_{\nu}\in{\mathbb{C}}$ and $w_{\nu}\in{\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap C_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})$, $\nu=1,\dots,m$, the functions ${\mathcal{F}}_{j}^{(\sigma)}={\mathcal{F}}_{j}^{(\sigma)}(\varepsilon,\tau,\{b_{\nu}\},\{w_{\nu}\})$, $\sigma=0,1$, belong to the space $C^{\,even}_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})$. If $w_{\nu}\in{\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap C^{\infty}_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})$, $\nu=1,\dots,m$, then ${\mathcal{F}}_{j}^{(\sigma)}\in C^{\infty,\,even}_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})$.
(2) If the numbers $\tau,b_{\nu}$ and the functions $w_{\nu}\in{\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap C_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})$, $\nu=1,\dots,m$, depend on a number $\varepsilon\in{\mathbb{C}}$, and functions ${\mathbb{C}}\ni\varepsilon\mapsto\tau\in{\mathbb{C}}$, ${\mathbb{C}}\ni\varepsilon\mapsto b_{\nu}\in{\mathbb{C}}$, and ${\mathbb{C}}\ni\varepsilon\mapsto w_{\nu}\in\bigl({\mathfrak{L}}({\mathbb{Y}}^{\prime})\,\cap\,C_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}}),\|\cdot\|_{C_{\Lambda(T)}}\bigr)$ are analytic, then functions ${\mathbb{C}}\ni\varepsilon\mapsto{\mathcal{F}}_{j}^{(\sigma)}(\varepsilon,\tau,\{b_{\nu}\},\{w_{\nu}\})\in\bigl(C^{\,even}_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}}),\|\cdot\|_{C_{\Lambda(T)}}\bigr)$ also are analytic.
(3) For any $R>0$ there exist numbers $C_{R}^{(\sigma)}>0$, $\sigma=0,1$, such that for all numbers $\varepsilon,\tau^{\prime},\tau^{\prime\prime},b_{\nu}^{\prime},b_{\nu}^{\prime\prime}\in\overline{U}_{R}$ and all functions $w_{\nu}^{\prime},w_{\nu}^{\prime\prime}\in{\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap C_{\Lambda}({\mathbb{R}};{\mathbb{C}})$, for which $\|w_{\nu}^{\prime}\|_{C_{\Lambda(T)}}\leqslant R$ and $\|w_{\nu}^{\prime\prime}\|_{C_{\Lambda(T)}}\leqslant R$, $\nu=1,\dots,m$, inequalities
$\sum\limits_{j}\|{\mathcal{F}}_{j}^{(\sigma)}(\varepsilon,\tau^{\prime},\{b_{\nu}^{\prime}\},\{w_{\nu}^{\prime}\})-{\mathcal{F}}_{j}^{(\sigma)}(\varepsilon,\tau^{\prime\prime},\{b_{\nu}^{\prime\prime}\},\{w_{\nu}^{\prime\prime}\})\|_{C_{\Lambda(T)}}$
$\leqslant\,C_{R}^{(\sigma)}\biggl(|\tau^{\prime}-\tau^{\prime\prime}|+\sum\limits_{\nu}|b_{\nu}^{\prime}-b_{\nu}^{\prime\prime}|+\sum\limits_{\nu}\|w_{\nu}^{\prime}-w_{\nu}^{\prime\prime}\|_{C_{\Lambda(T)}}\biggr)$ (4.8)
hold.
Lemma 11 implies that system of equations (4.7) is equivalent to the system
$w_{j}={\mathbb{W}}_{j}+\varepsilon{\mathcal{F}}_{j}^{(2)},\quad j=1,\dots,m,$ (4.9)
where
${\mathbb{W}}_{j}\doteq\frac{1}{2}\,\sum\limits_{\mu}\,\biggl(\,\sum\limits_{\nu}C^{(m)}_{\mu\nu}a^{2}_{\nu}\biggr)\widehat{S}_{j\mu}w^{2}_{*},\quad{\mathcal{F}}^{(2)}_{j}\doteq\sum\limits_{\mu}\widehat{S}_{j\mu}\widehat{P}_{\,{\mathbb{Y}}^{\prime}}{\mathcal{F}}^{(1)}_{\mu}$.
Therefore, if system of equations (4.7) is valid, then system of equations (4.6) is equivalent to the system
$\biggl(\omega^{2}b_{j}-B_{0}\sum\limits_{\mu}C^{(m)}_{j\mu}b_{\mu}-\tau\omega^{2}a_{j}\biggr)w_{*}=B_{0}\widehat{P}_{\,{\mathbb{Y}}}\widetilde{\mathbb{W}}_{j}+\varepsilon B_{0}\widehat{P}_{\,{\mathbb{Y}}}{\mathcal{F}}_{j}^{(3)},\quad j=1,\dots,m,$ (4.10)
where
$\widetilde{\mathbb{W}}_{j}\doteq\sum\limits_{\mu}C^{(m)}_{j\mu}a_{\mu}w_{*}{\mathbb{W}}_{\mu}+\frac{1}{6}\,\biggl(\,\sum\limits_{\mu}C^{(m)}_{j\mu}a^{3}_{\mu}\biggr)w^{3}_{*},\quad{\mathcal{F}}_{j}^{(3)}\doteq\sum\limits_{\mu}C^{(m)}_{j\mu}a_{\mu}w_{*}{\mathcal{F}}_{\mu}^{(2)}+{\mathcal{F}}^{(0)}_{j}$.
The functions ${\mathcal{F}}_{j}^{(\sigma)}={\mathcal{F}}_{j}^{(\sigma)}(\varepsilon,\tau,\{b_{\nu}\},\{w_{\nu}\})$, $\sigma=2,3$, as well as the functions ${\mathcal{F}}_{j}^{(\sigma)}={\mathcal{F}}_{j}^{(\sigma)}(\varepsilon,\tau,\{b_{\nu}\},\{w_{\nu}\})$, $\sigma=0,1$, satisfy properties (1), (2), and (3) (for the functions ${\mathcal{F}}_{j}^{(2)}$, property (2) is fulfilled by Lemma 12).
Since $\|w_{*}\|^{2}_{L^{2}_{\Lambda(T)}}=2T$, it follows from Lemma 10 that equations (4.10) hold for $\varepsilon\in{\mathbb{C}}$ iff
$\tau=\tau^{(0)}+\frac{\varepsilon}{2T}\,\sum\limits_{\mu}\tau_{\mu}(w_{*},{\mathcal{F}}^{(3)}_{\mu})_{L^{2}_{\Lambda(T)}},$ (4.11)
$b_{\nu}=b_{\nu}^{(0)}+\frac{\varepsilon}{2T}\,\sum\limits_{\mu}R_{\nu\mu}(w_{*},{\mathcal{F}}^{(3)}_{\mu})_{L^{2}_{\Lambda(T)}},\quad\nu=1,\dots,m,$ (4.12)
where
$\tau^{(0)}\doteq\frac{1}{2T}\,\sum\limits_{\mu}\tau_{\mu}(w_{*},\widetilde{\mathbb{W}}_{\mu})_{L^{2}_{\Lambda(T)}},\quad b_{\nu}^{(0)}\doteq\frac{1}{2T}\,\sum\limits_{\mu}R_{\nu\mu}(w_{*},\widetilde{\mathbb{W}}_{\mu})_{L^{2}_{\Lambda(T)}}.$
Denote $w^{(0)}_{\nu}\doteq{\mathbb{W}}_{\nu}$, $\nu=1,\dots,m$, and sequentially for $j=1,2,\dots$ let us define numbers $\tau^{(j)}=\tau^{(j)}(\varepsilon)$, $b_{\nu}^{(j)}=b_{\nu}^{(j)}(\varepsilon)$ and functions $w_{\nu}^{(j)}=w_{\nu}^{(j)}(\varepsilon;\cdot)$. If they are already found for some $j\in{\mathbb{Z}}_{+}$, then we set
$\tau^{(j+1)}=\tau^{(0)}+\frac{\varepsilon}{2T}\,\sum\limits_{\mu}\tau_{\mu}\big(w_{*},{\mathcal{F}}^{(3)}_{\mu}(\varepsilon,\tau^{(j)},\{b_{s}^{(j)}\},\{w^{(j)}_{s}\})\bigr)_{L^{2}_{\Lambda(T)}},$
$b_{\nu}^{(j+1)}=b_{\nu}^{(0)}+\frac{\varepsilon}{2T}\,\sum\limits_{\mu}R_{\nu\mu}\big(w_{*},{\mathcal{F}}^{(3)}_{\mu}(\varepsilon,\tau^{(j)},\{b_{s}^{(j)}\},\{w^{(j)}_{s}\})\bigr)_{L^{2}_{\Lambda(T)}},$
$w_{\nu}^{(j+1)}=w_{\nu}^{(0)}+\varepsilon{\mathcal{F}}^{(2)}_{\nu}(\varepsilon,\tau^{(j)},\{b_{s}^{(j)}\},\{w^{(j)}_{s}\})$.
Then $w_{\nu}^{(j)}(\varepsilon;\cdot)\in{\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap C^{\infty}_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})$ for all $\varepsilon\in{\mathbb{C}}$, and functions ${\mathbb{C}}\ni\varepsilon\mapsto\tau^{(j)}(\varepsilon)$, ${\mathbb{C}}\ni\varepsilon\mapsto b_{\nu}^{(j)}(\varepsilon)$, ${\mathbb{C}}\ni\varepsilon\mapsto w_{\nu}^{(j)}(\varepsilon;\cdot)\in\bigl({\mathfrak{L}}({\mathbb{Y}}^{\prime})\cap C_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}}),\|\cdot\|_{C_{\Lambda(T)}}\bigr)$ are analytic (and $\sum\limits_{\nu}\overline{b}^{(j)}_{\nu}(\varepsilon)a_{\nu}=0$), $\nu=1,\dots,m$, $j\in{\mathbb{N}}$. If $\varepsilon\in{\mathbb{R}}$, then $\tau^{(j)}(\varepsilon),b_{\nu}^{(j)}(\varepsilon)\in{\mathbb{R}}$ and $w^{(j)}_{\nu}(\varepsilon;\cdot)\in C^{\infty}_{\Lambda(T)}({\mathbb{R}};{\mathbb{R}})$.
[…]
$\bigl|\tau^{(j+1)}(\varepsilon)-\tau^{(j)}(\varepsilon)\bigr|+\sum\limits_{\nu}\,\bigl|b_{\nu}^{(j+1)}(\varepsilon)-b_{\nu}^{(j)}(\varepsilon)\bigr|+\sum\limits_{\nu}\,\bigl\|w_{\nu}^{(j+1)}(\varepsilon;\cdot)-w_{\nu}^{(j)}(\varepsilon:\cdot)\bigr\|_{C_{\Lambda(T)}}$
$\leqslant\,\frac{1}{2}\,\biggl(\bigl|\tau^{(j)}(\varepsilon)-\tau^{(j-1)}(\varepsilon)\bigr|+\sum\limits_{\nu}\,\bigl|b_{\nu}^{(j)}(\varepsilon)-b_{\nu}^{(j-1)}(\varepsilon)\bigr|+\sum\limits_{\nu}\,\bigl\|w_{\nu}^{(j)}(\varepsilon;\cdot)-w_{\nu}^{(j-1)}(\varepsilon:\cdot)\bigr\|_{C_{\Lambda(T)}}\biggr).$
Hence for any $\varepsilon\in\overline{U}_{r}$, there exist numbers $\tau(\varepsilon)$, $b_{\nu}(\varepsilon)\in{\mathbb{C}}$ and functions $w_{\nu}(\varepsilon;\cdot)\in{\mathfrak{L}}({\mathbb{Y}}^{\prime})\,\cap\,C^{\infty}_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})$ such that $\tau^{(j)}(\varepsilon)\to\tau(\varepsilon)$, $b_{\nu}^{(j)}(\varepsilon)\to b_{\nu}(\varepsilon)$, and $\|w_{\nu}^{(j)}(\varepsilon;\cdot)-w_{\nu}(\varepsilon;\cdot)\|_{C_{\Lambda(T)}}\to 0$ as $j\to+\infty$. Because the convergence is uniform in $\varepsilon\in\overline{U}_{r}$, it follows that the functions $U_{r}\ni\varepsilon\mapsto\tau(\varepsilon)$, $U_{r}\ni\varepsilon\mapsto b_{\nu}(\varepsilon)$, and $U_{r}\ni\varepsilon\mapsto w_{\nu}(\varepsilon;\cdot)\in\bigl({\mathfrak{L}}({\mathbb{Y}}^{\prime})\,\cap\,C^{\infty}_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}}),\|\cdot\|_{C_{\Lambda(T)}}\bigr)$ are analytic. For all $\varepsilon\in\overline{U}_{r}$, the numbers $\tau(\varepsilon)$, $b_{j}(\varepsilon)$ and the functions $w_{j}(\varepsilon;\cdot)$ satisfy equations (4.11), (4.12), and (4.9), therefore $w_{j}(\varepsilon;\cdot)\in H^{2}_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})\subset C_{\Lambda(T)}({\mathbb{R}};{\mathbb{C}})$ and functions (4.5) are solutions of system (1.11). In addition, the number $r>0$ can be chosen sufficiently small such that the inequality ${\varepsilon}^{2}|\tau(\varepsilon)|<1$ holds for all $\varepsilon\in\overline{U}_{r}$. For real numbers $\varepsilon\in(-r,r)$, the numbers $\tau(\varepsilon)$, $b_{j}(\varepsilon)$ also are real and the functions $u_{j}(\varepsilon;\cdot)$ belong to $H^{2}_{\Lambda(T)}({\mathbb{R}};{\mathbb{R}})$. Since under the condition $u_{j}(\varepsilon;\cdot)\in H^{2n}_{\Lambda(T)}({\mathbb{R}};{\mathbb{R}})$, $n\in{\mathbb{N}}$, $j=1,\dots,m$, it follows from (1.11) that $u_{j}(\varepsilon;\cdot)\in H^{2(n+1)}_{\Lambda(T)}({\mathbb{R}};{\mathbb{R}})$, we have
$u_{j}(\varepsilon;\cdot)\in\bigcap\limits_{n\,\in\,{\mathbb{N}}}H^{2n}_{\Lambda(T)}({\mathbb{R}};{\mathbb{R}})=C^{\infty}_{\Lambda(T)}({\mathbb{R}};{\mathbb{R}}).$
The proof of Theorem 7 is complete.